The speed of a boat in still water is 5 miles per hour. If the boat travels 3 miles downstream in the same amount of time it takes to travel 1.5 miles upstream, what is the speed of the current?
Answer by mananth(16946)

boat speed = 5 mph
current speed = x mph

Speed upstream= 5 - x
Speed downstream= 5 + x

Distance upstream= 1.5 miles
Distance downstream= 3 miles

t=d/r time upstream = time downstream
1.5 / ( 5 - x )= 3 / ( 5 + x )
1.5 ( 5 + x )= 3 ( 5 -x )
7.5 + 1.5 x = 15 - 3 x
3 x -1.5 x = 7.5
1.5 x = 7.5
/ 1.5
x = 5 mph boat speed